Can you please tell me if I have this right? If not, can you please explain... Thanks sooo much!
Assume a class has 30 members. How many committees of three people can be chosen?
I have:
30*29/3=290

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

Assume a class has 30 members. How many committees of three people can be chosen?
----------------
A committee is a group of people so this is a combination problem.
----------------
30C3 = 30!/[(30-3)!*3!]
---
= (30*29*28)/(1*2*3)
= 4060 committees
========================
Cheers,
Stan H.

Answer by Alan3354(69443)

Assume a class has 30 members. How many committees of three people can be chosen?
------------
30/3 = 10 committees.
---------------------
If you meant in how many different ways can a committee of 3 be chosen:
The 1 choice is 1 of 30.
Then 1 of 29, then 28.
--> 30*29*28 = 24360.
But, a committee of A, B and C is the same as C, A and B, so it's necessary to divide by 6 (3*2*1 = 6)
--------------------------
24360/6 = 4060 different combinations.